OpenStudy (anonymous):

Find the volume of a cube with the given side length. 10 units A. 10,000 cubic units B. 1000 cubic units C. 100 cubic units D. 30 cubic units

OpenStudy (jango_in_dtown):

If a is the length of a side of a cube then its volume its a^3

OpenStudy (jango_in_dtown):

here a=10units so a^3=10*10*10 cubic units =1000cubic units